a father is 27 years older than his son.the sum of their age is 55 years.find the age of the son?
Answer by Finavon(81)

S and F: ages of son and father
S+27=F (1)
S+F=55 (2)
Substitute for F from (1)
(2) S+S+27=55
so 2S = 28
Son is 14
Father 41